1. Blood is maintained at a pH of 7.4 by the primary buffers in the plasma and secondary buffers in the erythrocytes. The plasma contains carbonic acid/bicarbonate as one of its primary buffers. The pK for this buffer system is 6.3

H2CO3 <=> H+ + HCO3-

What is the ratio of bicarbonate to carbonic acid in normal blood plasma?
Answer #1

Solution:

According to Hendersen-Hasselbalch equation for buffer solution,

pH = pKa + log [salt] / [Acid]

For bicarbonate ( HCO3-) and carbonic acid (H2CO3) buffer above equation becomes as,

pH = pka + log [HCO3-] / [H2CO3]

7.4 = 6.3 + log [HCO3-] / [H2CO3]

log [HCO3-] / [H2CO3] = 7.4 - 6.3 = 1.1

[HCO3-] / [H2CO3] = antilog 1.1 = 12.589 = 12.59

Hence, ratio of bicarbonate to carbonic acid = 12.59